# Bayesian Analysis of High Dimensional Vector Error Correction Model

### 28 Dec 2023 v1

@@ 2.3 Algorithms
In general, the algorithms employ EM techniques to numerically approximate the estimators $(\hat{R}^{BE},\hat{\sigma}^{BE},\hat{\theta}^{ABE})$ as per Equation 11 and Equation 15. Write these as $(\hat{R},\hat{\sigma},\hat{\theta})$ and the approximated rank $\hat{r}:=rank(\hat{R})$. The two algorithms differ by their approaches in $\lambda_{0}$ searching, in that algorithm 1 assumes the same value of $\lambda_{0,j}=\lambda_{0}\forall j$ whereas algorithm 2 allows some degree of different increments of $\lambda_{0,j}$ and associated termination. Such difference is propagated through a uniform sampling on non-sparse coordinates.
@@ 2.3.1 Algorithm 1
Initialisers $\theta(0),\beta(0),\sigma_{j}^{2}(0)$;
@@
Parametric specifications for prior on $\theta_{i}$: $a_{i}$, $b_{i}$
@@ 3             Initialise $\theta_{j}(0),\beta_{j}(0),\sigma_{j}^{2}(0)$, then for $k\in[K]$ do
$\theta_{j}(k)=$ $\frac{a-1+\sum_{i\in[p]}p_{\theta_{j}(k-1)}^{*}(\beta_{i,j}(k))}{a+b+p-2}$
@@ 9            Collect $\hat{\beta_{j}}=\beta_{j}(k)$
2.3.2 Algorithm 2
Input: Data A,B ;
@@ 2.3.2 Algorithm 2
Initialisers $\theta(0),\hat{R}(0),\sigma^{2}(0)$;
@@
Parametric specifications for prior on $\theta_{i}$: $a_{i}$, $b_{i}$
Seed for uniform sampling
@@
2       Use algorithm 1 to update $\lambda_{0}$ with $\lambda_{0}+\Delta_{\lambda}$ and EM update on $(\beta,\theta,\sigma^{2})$ Collect $\hat{r}(\lambda_{0})=rank\left(\hat{R}(\lambda_{0})\right)$
@@
$\theta_{j}(k)=$ $\frac{a-1+\sum_{i\in[p]}p_{\theta_{j}(k-1)}^{*}(\beta_{i,j}(k),j)}{a+b+p-2}$
@@ 3.1 Outline of Simulated Examples
with initialisation $Y_{0}=0$. If we write Equation 26 in the form of Equation 2, then
@@
Output: Estimated ranks $\hat{r}(i)$ for $i\in[S]$
@@ 3.2.1 Algorithm 1
Figure 2: Estimated ranks over iterations
@@ 3.2.2 Algorithm 2
For algorithm 2, we demonstrate the advancement from algorithm 1 in the following two ways: better performance (with reasons) under the high-dimensional ($p=30,T=300$) regime, and the ability to handle even higher-dimensional ($p=100,T=500$) regime. In both regimes, we experiment $r\in\{1,5\}$.
@@
Figure 5: Histograms of estimated ranks over samples for $p=100,T=500$
Figure 6: Estimated ranks over iterations
@@ 4.1 Portfolio construction
After obtaining $\hat{R}$ and $\hat{r}$, the key idea is to use $\hat{R}$ to reconstruct $\hat{\Pi}$ and hence obtain $\hat{r}$ number of co-integrating vectors. These vectors, upon normalisation, yield exactly a portfolio which we may analyse, assuming a single dollar and zero-cost implementation on a long-short portfolio on the relevant universe of assets.
Let $\hat{\Pi}$ be the recovered $\Pi$ where $\hat{\Pi}=(\tilde{S}\hat{R})^{\intercal}$. Let $\hat{\Pi}_{i}$ be a non-zero row vector of $\hat{\Pi}$. A portfolio is therefore constructed with value at time $t$ being $\alpha_{i}Y_{t}$, where $\alpha_{i}:=\frac{\hat{\Pi_{i}}}{||\hat{\Pi_{i}}||_{1}}$.
@@ 4.2 Data and Preparation
We obtain dailyt stock prices of all individual stocks in SPX 500 as of end of September 2023. The price time series start from January 2022 to September 2023 (this gives 374 observations, though it also means $T<p$), and we use from the January 2022 to June 2023 for in-sample training, and July to September 2023 for out-of-sample testing. We use ADF test to check for I(1) non-stationarity. Out of the 500 stocks, 405 stocks are I(1). We are henceforth in a regime of $T=374$ and $p=405$.
The only pre-processing to be done is to compute for cumulative return in relation to the start of the dataset, that is, to have $\frac{Y_{t,j}}{Y_{1,j}}$ as the time series for investigation.
@@ 4.3.1 Inference on estimated rank in algorithm 2
In Figure 7, we plot the estimated ranks over iterations on the left, and the histogram of $\{\hat{r}(s):s\in S\}$ on the right. An average rank of around 1.6 is being estimated.
Figure 7: Path of estimated ranks over stages of $\lambda_{0,j}$ increments (left); Histogram of estimated ranks over different seeds (right)
@@ 4.3.2 General training and testing performance
In Figure 8, we plot histograms of the training and testing volatilities of all such portfolios over all seeds. The values are much lower than the underlining index in both training and testing (see Table 2 for the index and equally-weighted portfolio volatility in the corresponding periods).
@@
Figure 8: Portfolio volatility in training and testing over different seeds
@@ 5.1 Discussions
A number of future works are thought to be helpful based on the current investigation: we may investigate a even larger number of features, either in the numerical sense (e.g. $p>1000$) or in the asymptotic sense of $p=O(T)$. The asymptotic inference may also be drawn towards approximate correctness, such as $\hat{r}/r\to 1$ with more generic limits such as $q,r\to\infty$. This then leads to space for investigating even higher dimensional cointegration in applied fields, such as portfolio construction and adjustments.
@@ 5.2 Contribution
There are mainly three forces of contributions being brought from this paper. First, similar to [5] and [2], we extend the ”Traditional identifiability constraint” from $p^{2}<T$ to $pr\leq T$, in the sense that through penalisation, one does not necessarily need more observations than unknowns to estimate the cointegration matrix and thereafter the number of cointegration relationships in the system. Instead, through SSL, we proposed nuanced methodologies to estimate $r$ and also showed some level of asymptotic convergence in a Bayesian setting.
Secondly, we used various simulation regimes to test our proposed algorithms. This furthers the SSL executions from optimally finding a $p$ dimensional sparse vector to finding a $p\times p$ dimensional sparse matrix and determining its true rank. Particularly, the termination of $\lambda_{0}$ increment varies substantially from existing Bayesian literature, as we quantify further the termination criteria instead of a generic guidance on convergence-termination. This contributes to the implementation of SSL and furthered the agenda to use Bayesian modelling as a ”pragmatic frequentist” on penalisation.

### 12 Mar 2024 v2

@@ 2.4 Algorithms
In general, the algorithms employ EM techniques (see [1] for further details on the EM techniques) to numerically approximate the estimators $(\hat{R}^{BE},\hat{\sigma}^{BE})$ as per Equation 14. Write these as $(\hat{R},\hat{\sigma})$ and the approximated rank $\hat{r}:=rank(\hat{R})$. The two algorithms differ by their approaches in $\lambda_{0}$ searching, in that algorithm 1 assumes the same value of $\lambda_{0,j}=\lambda_{0}\forall j$ whereas algorithm 2 allows some degree of different increments of $\lambda_{0,j}$ and associated termination. Such difference is propagated through a uniform sampling on non-sparse coordinates.
@@ 2.4.2 Algorithm 2
Initialisers $\theta(0),\beta(0),(\sigma_{j}^{2}(0))_{j\in[p]}$;
@@
Parametric specifications for prior on $\theta_{j}$: $a_{j}$, $b_{j}$
@@ 3             Initialise $\theta_{j}(0),\beta_{j}(0),\sigma_{j}^{2}(0)$, then for $k\in[K]$ do
$\theta_{j}(k)=$ $\frac{a_{j}-1+\sum_{i\in[p]}p_{\theta_{j}(k-1)}^{*}(\beta_{i,j}(k))}{a_{j}+b_{j}+p-2}$
@@ 9            Collect $\hat{\beta_{j}}=\beta_{j}(k)$
Input: Data A,B ; Seed for uniform sampling;
@@
Initialisers $\theta(0),\beta(0),(\sigma_{j}^{2}(0))_{j\in[p]}$;
@@
Parametric specifications for prior on $\theta_{j}$: $a_{j}$, $b_{j}$
@@
2       Use algorithm 1 to update $\lambda_{0}$ with vectorised $\lambda_{0}+\Delta_{\lambda}$ and EM update on $(\beta,\theta,\sigma^{2})$ Collect $\hat{r}(\lambda_{0})=rank\left(\hat{R}(\lambda_{0})\right)$
@@
$\theta_{j}(k)=$ $\frac{a_{j}-1+\sum_{i\in[p]}p_{\theta_{j}(k-1)}^{*}(\beta_{i,j}(k),j)}{a_{j}+b_{j}+p-2}$
@@ 3.1 Outline of Simulated Examples
with initialisation $Y_{0}=0$. $J(0,r)$ refers to the Jordan Block of dimension $r$ with diagonal taking value 0. If we write Equation 29 in the form of Equation 2, then
@@
Output: Estimated ranks $\hat{r}(i)$ for $i\in I$ with $|I|=S$
@@ 3.2.1 Algorithm 1
Figure 2: Estimated ranks over iterations of $\lambda_{0,j}$ increments
@@ 3.2.2 Algorithm 2
For algorithm 2, we demonstrate the advancement from algorithm 1 in the following two ways: better performance under the high-dimensional ($p=30,T=300$) regime, and the ability to handle even higher-dimensional ($p=100,T=500$) regime. In both regimes, we experiment $r\in\{1,5\}$.
@@
Figure 4: Estimated ranks over iterations of $\lambda_{0,j}$ increments
Figure 5: Histograms of estimated ranks over samples for $p=100,T=500$
@@
Figure 6: Estimated ranks over iterations of $\lambda_{0,j}$ increments
@@ 4.1 Portfolio construction
Recall in stock portfolio setting that $Y_{t}$ is a $p$-dimensional vector of stock price at time $t$, encompassing $p$ number of stocks. After obtaining $\hat{R}$ and $\hat{r}$, the key idea is to use $\hat{R}$ to reconstruct $\hat{\Pi}$ and hence obtain $\hat{r}$ number of co-integrating vectors. These vectors, upon normalisation, yield exactly a portfolio which we may analyse, assuming a single dollar and zero-cost implementation on a long-short portfolio on the relevant universe of assets.
Let $\hat{\Pi}$ be the recovered $\Pi$ where $\hat{\Pi}=(\tilde{S}\hat{R})^{\intercal}$. Let $\hat{\Pi}_{i}$ be a non-zero row vector of $\hat{\Pi}$. A portfolio is therefore constructed with value at time $t$ being $\alpha_{i}Y_{t}$, where $\alpha_{i}:=\frac{\hat{\Pi_{i}}}{||\hat{\Pi_{i}}||_{1}}$. The estimated rank $\hat{r}$ would henceforth be the number of such portfolios, as $\alpha_{1},...,\alpha_{\hat{r}}$ are linearly independent cointegration vectors.
@@ 4.2 Data and Preparation
We obtain daily stock prices of all individual stocks in SPX 500 as of end of September 2023. The price time series start from January 2022 to September 2023 (this gives 374 observations, though it also means $T<p$), and we use from the January 2022 to June 2023 for in-sample training, and July to September 2023 for out-of-sample testing. We use ADF test to check for I(1) non-stationarity. Out of the 500 stocks, 405 stocks are I(1). We are henceforth in a regime of $T=374$ and $p=405$.
The only pre-processing is to compute for cumulative return in relation to the start of the dataset, that is, to have $\frac{Y(t,j)}{Y(1,j)}$ as the time series for investigation.
@@ 4.3.1 Inference on estimated rank in algorithm 2
In Figure 7, we plot the estimated ranks over iterations on the left, and the histogram of $\{\hat{r}(s):s\in S\}$ on the right. An average rank of 1.7 is being estimated, as highlighted in the red verical line of the histogram.
Figure 7: Path of estimated ranks over iterations of $\lambda_{0,j}$ increments (left); Histogram of estimated ranks over different seeds (right)
@@ 4.3.2 General training and testing performance
In Figure 8, we plot histograms of the training and testing volatilities of all such portfolios over all seeds, which gives 170 portfolios as the average rank is 1.7. The values of training and testing volatilities are much lower than the underlining index in both training and testing (see Table 2 for the index and equally-weighted portfolio volatility in the corresponding periods).
@@
Figure 8: Portfolio volatility in training (left) and testing (right) over the 170 implied portfolios
@@ 5.1 Discussions
A number of future works are thought to be helpful based on the current investigation: we may investigate a even larger number of features, either in the numerical sense (e.g. $p>1000$) or in the asymptotic sense of $p=O(T)$. The asymptotic inference may also be drawn towards approximate correctness, such as $\hat{r}/r\to 1$ with more generic limits such as $q,r\to\infty$. This then leads to space for investigating even higher dimensional cointegration in applied fields, such as portfolio construction and adjustments. It is also worth extending the SSL-VECM Bayesian model on covariance matrices, which may play a role in some cointegrated relationships. Though, such covariance may be sparse, adding another level of difficulty for the model estimation.
@@ 5.2 Contribution
There are mainly three forces of contributions being brought from this paper. First, similar to [7] and [4], we extend the ‘traditional identifiability constraint’ from $p^{2}<T$ to $pr\leq T$, in the sense that through penalisation, one does not necessarily need more observations than unknowns to estimate the cointegration matrix and thereafter the number of cointegration relationships in the system. Instead, through SSL, we proposed nuanced methodologies to estimate $r$ and also showed some level of asymptotic convergence in a Bayesian setting.
Secondly, we used various simulation regimes to test our proposed algorithms. This furthers the SSL executions from optimally finding a $p$ dimensional sparse vector to finding a $p\times p$ dimensional sparse matrix and determining its true rank. Particularly, the termination of $\lambda_{0}$ increment varies substantially from existing Bayesian literature, as we quantify further the termination criteria instead of a generic guidance on convergence-termination. This contributes to the implementation of SSL and furthered the agenda to use Bayesian modelling as a ‘pragmatic frequentist’ on penalisation.
@@ A.1 Data processing
A.2 Derivation to Equation 9
First, we note $\pi(\phi|D,D_{0})\propto\pi(D|\phi,D_{0})\pi(\phi)$ where $\pi(D|\phi,D_{0})$ is the likelihood conditional on some initialisation $D_{0}$ and $\pi(\phi)$ is the prior of $\phi$. $\propto$ refers to proportional up to some constants not involving $\phi$.
Now, we can write
$\pi(\phi|D,D_{0})$ $\propto\pi(D|R,\sigma^{2},D_{0})\pi(\sigma^{2})\pi(R)$ (32)
$=\left(\prod_{t\in[T]}\pi(\tilde{A}_{t}|\tilde{B}_{t},R,\sigma^{2})\right)\pi(\sigma^{2})\left(\prod_{j\in[p]}\sum_{\gamma_{j}\in\{0,1\}}\pi(R(\cdot,j)|\gamma_{j})\pi(\gamma_{j}|\theta_{j})\right)$ (33)
We further observe, for any $t\in[T]$, $j\in[p]$,
$\pi(\tilde{A}_{t}|\tilde{B}_{t},R,\sigma^{2})$ $\propto\sigma^{-p}\exp\left(-\frac{||e_{t}(R)||_{2}^{2}}{2\sigma^{2}}\right)$ (34)
$\sum_{\gamma_{j}\in\{0,1\}}\pi(R(\cdot,j)|\gamma_{j})\pi(\gamma_{j}|\theta_{j})$ $=(1-\theta_{j})\psi_{0,j}(R(\cdot,j))+\theta_{j}\psi_{1}(R(\cdot,j))$ (35)
$\propto\exp(-\lambda_{1}||R(\cdot,j)||_{1})\times\prod_{i\in[p]}\frac{1}{p_{j}^{*}(R(i,j),\theta_{j})}$ (36)
Henceforth combining everything, we have
$\pi(\phi|D,D_{0})\propto\left(\prod_{t\in[T]}\exp\left(-\frac{||e_{t}(R)||_{2}^{2}}{2\sigma^{2}}\right)\right)\sigma^{-(pT+2)}\prod_{j\in[p]}\left(\exp(-\lambda_{1}||R(\cdot,j)||_{1})\times\prod_{i\in[p]}\frac{1}{p_{j}^{*}(R(i,j),\theta_{j})}\right)$
Then we take logarithm and obtain
$(\hat{R}^{BE},\hat{\sigma}^{BE})=\operatorname*{argmin}_{R\in\mathbb{R}^{p\times p},\sigma\in\mathbb{R}_{>0}}-\log(\pi(\phi|D,D_{0}))$ (37)
to the desired Equation 9.
A.3 Derivation to Equation 14
Firstly, breaking down we obtain
[…]
Similar to Appendix A.2, we proceed to observe, for any $t\in[T]$, $j\in[p]$,
$\pi(\tilde{A}_{t,j}|\tilde{B}_{t},R,\sigma_{j}^{2})$ $\propto\sigma_{j}^{-1}\exp\left(-\frac{(e_{t}(R))_{j}^{2}}{2\sigma_{j}^{2}}\right)$ (39)
$=\sigma_{j}^{-1}\exp(-\epsilon(j)^{2}_{t})$ (40)
$\pi(R(\cdot,j)|\theta_{j})\pi(\theta_{j})$ $=((1-\theta_{j})\psi_{0,j}(R(\cdot,j))+\theta_{j}\psi_{1}(R(\cdot,j)))\pi(\theta_{j})$ (41)
$=\exp(-\lambda_{1}||R(\cdot,j)||_{1})\times\prod_{i\in[p]}\frac{\theta_{j}\pi(\theta_{j})}{p_{j}^{*}(R(i,j),\theta_{j})}$ (42)
Therefore
$\pi(\phi|D,D_{0})$ $\propto\left(\prod_{j\in[p]}\exp(-||\epsilon(j)||_{2}^{2})\sigma_{j}^{-(T+2)}\right)\left(\prod_{j\in[p]}\left(\exp(-\lambda_{1}||R(\cdot,j)||_{1})\times\int_{(0,1)}\prod_{i\in[p]}\frac{\theta_{j}\pi(\theta_{j})}{p_{j}^{*}(R(i,j),\theta_{j})}d\theta_{j}\right)\right)$ (43)
Then we take logarithm and obtain
$(\hat{R}^{BE},(\hat{\sigma}_{j}^{BE})_{j\in[p]})=\operatorname*{argmin}_{R\in\mathbb{R}^{p\times p},({\sigma}_{j}^{BE})_{j\in[p]}\in\mathbb{R}^{p}_{>0}}-\log(\pi(\phi|D,D_{0}))$ (44)
to the desired Equation 14.